### v2

@@ 1.4 Outline of the Paper
In Section 2, we present the formal model of shill bidding in auctions. Section 3 studies Dutch auctions and their benefits with respect to disincentivizing shill bidding. Section 4 explores which formats are both weakly shill-proofness and ex-post equilibrium for real bidders. Section 5 presents our trilemma result for single-action auctions. And in Section 6, we conclude the paper by discussing extensions.
@@ 2 Model
In this paper, we consider extensive-form, single item auctions. An extensive-form game $G$ is a tuple of possible histories $H$, and, for each history $h\in H$, functions mapping $h$ to: (i) a player taking an action, $P(h)$; (ii) a set of possible actions, $A(h)$; (iii) an information set, [^16] The information sets induce a partition over all possible histories. Players’ actions can condition only on $\mathcal{I}(h)$, not on $h$. $\mathcal{I}(h)$; and (iv) the most recent action taken, $\mathcal{A}(h)$. As further notation, we denote the starting history of the game by $h_{\emptyset}$ and the set of terminal histories as $Z$; we say $h^{\prime}\prec h$ if $h^{\prime}$ precedes $h$, i.e., there exists a sequence of actions that lead from $h^{\prime}$ to $h$.
We restrict attention to single item auctions, which means that for every terminal history $z\in Z$, we can associate an allocation and transfer vector: $z=(x,t)$, with $\sum_{i=1}^{N}x_{i}\leq 1$ and $x_{i}\in\left\{0,1\right\}$ for all $i$. As abuse of notation, we will use $x(z),t(z)$ to mean the vectors $(x,t)$ associated with the terminal history $z$. We also assume perfect recall and finite depth. (Definition O.1 in the Online Appendix gives a formal, thorough, and standard definition of extensive form games.) Note that bidders’ values for the item are not built into the extensive-form game $G$. Instead, a strategy is a function of both the information set of a bidder and her value for the item. Like most papers in the extensive-form auction literature, we study games with a finite type space because defining auctions with a continuum of types requires defining a general class of continuous-time games. To the authors’ knowledge, there is no theory of continuous time games that rivals the generality and flexibility of extensive-form games.
@@ 2.1 Bidders – Real and Shill
In the auction, there is a set of potential bidders $B$, with $\left|B\right|=N$, who might participate. Of these potential bidders, a set of real bidders $R$ actually participates. Each bidder $i\in B$ has an independent probability $p$ of participating, $\mathds{P}\left[i\in R\right]=p$ [^17] This randomness plays little role in our analysis—we impose it only so that the overarching structure of our model has bidders with ex-ante, symmetric, independent private values. See also Footnote 10. The other bidders, $S=B\backslash R$, are shill bidders whose incentives are completely aligned with the seller/auctioneer’s, i.e., their utility is defined by the sum of real bidders’ transfers: for $i\in S,u_{i}(z)=\sum_{j\in R}t_{j}(z)$. Importantly, the auction $G$ cannot directly condition on the realization of $R$, i.e., the shill bidders are indistinguishable from real bidders during the auction. Each real bidder $i\in R$ has value $v_{i}$ for the item being sold where $v_{i}\sim F$ independently for each $i$. Each real bidder has quasi-linear utility: for $i\in R$, $u_{i}(z)=x_{i}(z)v_{i}-t_{i}(z).$ We assume $F$ is discrete, with support $\mathcal{V}$ consisting of the ordered atoms $0=v^{1}<v^{2}<\ldots<v^{M}$, and we define $f(v^{k})=\mathds{P}_{w\sim F}\left[w=v^{k}\right]$ to be the pmf of the distribution. As notation, for each shill bidder $i\in S$, we assign $v_{i}=0$ and let $v=(v_{1},\ldots,v_{N})$. The choice of values for shill bidders does not affect their incentives, and by supposing that their values are 00, we can define efficiency and optimality (revenue maximizing) in terms of only $v$ instead of $v$ and $R$ [^18] Here, we assume that the seller has 00 value for the item. Furthermore, when considering optimal auctions, we naturally assume the seller only cares about raising revenue from real bidders. Observe that given how $v$ is generated, we are in the standard, symmetric, single-item independent private values (IPV) setting.
@@
Real bidders have no information about who else is a real bidder. [^19] Unlike the assumption that $N$ is fixed, this assumption is an economically substantive one: The only way for shill bidders to manipulate the outcome of the auction is for shill bidders to take actions in the auction. If the bidders instead knew who were the real bidders, shill bidders could have an incentive to appear as real bidders in order to increase perceived competition. We do not explicitly consider the possibility that real bidders may update about which bidders may be shills over the course of the auction because under our shill-proofness conditions, shill bidders will never make nontrivial bids in equilibrium. We assume that shill bidders know the set of shill bidders and observe all previous actions taken. Formally, for any history $h$, if $P(h)\in S$, then $\mathcal{I}(h)=\left\{h\right\}$. This assumption rules out games with simultaneity (including static games) from the perspective of the shill bidders, but not real bidders. [^20] In extensive form games, simultaneity is modeled as the information set of a player having multiple elements. Without cryptography or other unmodeled technologies, we view it as reasonable to assume that while the auction may appear simultaneous to the real bidders, actions are taking place sequentially and the seller can observe those actions.
Our equilibrium concept is pure-strategy Perfect Bayesian Equilibrium; a formal definition of the auction equilibrium $(G,\sigma)$ can be found in Definition A.1. We write $\sigma(v;R)$ for the strategy profile when the value profile is $v$ and the realized set of real bidders is $R$.
@@ 2.2 Auction Environment
Throughout the paper, we focus on auction equilibria that are ex-post monotone and individually rational: The auction equilibrium $(G,\sigma)$ is monotone if, for all $i$, $j$, $v_{-j}$, and $v_{j}>v_{j}^{\prime}$, $\left[t_{i}\left(\sigma(v;B)\right)>0\implies t_{i}\left(\sigma(v;B)\right)\geq t_{i}\left(\sigma(v_{j}^{\prime},v_{-j};B)\right)\right],$ and is individually rational (IR) if, for all $v$ and $i\in B$, $x_{i}(\sigma(v;B))v_{i}-t_{i}(\sigma(v;B))\geq 0.$
@@
Because the value distribution is discrete, we must consider what to do if multiple bidders have the same (highest) value. We assume throughout the paper the notion of orderliness introduced by Akbarpour and Li (2020): there exists a fixed priority order—independent of values—over which bidder wins an item if there is a tie. [^21] For example, if ties are broken lexicographically, then the auction is orderly.
@@
    for all $v,i,j;v_{i}>v_{j}\implies(v_{i},i)\rhd(v_{j},j)$; and
@@
    for all $i,j$, if there exists $m$ such that $(v^{m},i)\rhd(v^{m},j)$, then for all $k$, $(v^{k},i)\rhd(v^{k},j)$.
In order to give the optimal auction a well-behaved allocation rule, we suppose the value distribution is regular (with bidders’ private values drawn i.i.d.). In Section 5, we relax the regularity assumption. We take our definition of a discrete regular distribution from Elkind (2007):
Definition 2.2.
A distribution $F$ is regular, if for all $k$, the virtual value $\varphi^{k}=v^{k}-(v^{k+1}-v^{k})\frac{1-F(v^{k})}{f(v^{k})}$ is non-decreasing.
[…]
2.3 Shill-Proofness
Next, we define our key shill-proofness desiderata. We are interested in auction equilibria in which shill bidders do not shill. Formally, this corresponds to requiring that shill bidders always act like real bidders who have value 00 for the item—since real bidders who have value 00 will never enter non-trivial bids in equilibrium, requiring shill bidders to have the same actions in equilibrium in effect means that shilling does not occur.
Definition 2.3.
An auction equilibrium $(G,\sigma)$ is weakly shill-proof if $\sigma$ is invariant to the realization of $S$, i.e., for all $v$ and $S,S^{\prime}\subseteq\left\{i:v_{i}=0\right\}$: $\sigma\left(v;B\backslash S\right)=\sigma\left(v;B\backslash S^{\prime}\right)$.
We can also strengthen our notion of shill-proofness from not shilling being an equilibrium strategy to being an ex-post strategy:
Definition 2.4.
An auction equilibrium $(G,\sigma)$ is strongly shill-proof if it is weakly shill-proof and a ex-post strategy profile for shill bidders, i.e., for all $\sigma^{\prime}$, $S$, and $v_{-S}$,
$\sum_{j\in R}t_{j}(\sigma(0,v_{-S};R))\geq\sum_{j\in R}t_{j}\left(\sigma_{S}^{\prime},\sigma_{-S}(0,v_{-S};R)\right).$
Strong shill-proofness is obviously preferable to weak shill-proofness (all else equal), especially if there are concerns about a seller somehow acquiring information about real bidders’ valuations beyond what is encoded by the prior. As we’ll see, however, the design space of weak shill-proof auctions is meaningfully larger than that of strong shill-proof auctions.
2.4 Revelation Principle
In order to make progress in understanding shill-proof auction formats, the following revelation principle will be helpful: for every auction equilibrium $\left(G,\sigma\right)$, there exists a direct auction that can be summarized by a direct allocation rule $\tilde{x}$, a direct transfer rule $\tilde{t}$, a menu rule $\mu$, and a starting player $\xi_{0}$. [^23] This revelation principle is similar to those found in, for example, Ashlagi and Gonczarowski (2018); Mackenzie (2020); Mackenzie and Zhou (2022); Pycia and Troyan (2023). The first input to the menu rule $\mu$ is a set $V$ of valuation profiles of the form $V_{1}\times V_{2}\times\cdots\times V_{N}$ with $V_{i}\subseteq\mathcal{V}$ for all $i$—intuitively, the valuation profiles that are, in equilibrium, consistent with a particular history. The second input is a player $\xi$ who is to move next. The output of the rule is a collection $\left\{\left(W_{\ell},\vec{\xi}_{\ell}\right)\right\}_{\ell\in\left\{1,\ldots,L\right\}}$, where the $W_{\ell}$’s are a partition of $V_{\xi}$ (from which player $\xi$ will choose one, according to her valuation, the equilibrium strategy $\sigma$ determines the partition) and $\vec{\xi}_{\ell}$ indicates the next player to move should player $\xi$ choose $W_{\ell}$. Under $\sigma$, the player $\xi$ will always select the partition $W_{\ell}$ such that $v_{\xi}\in W_{\ell}$. If $\vec{\xi}_{\ell}=\emptyset$, then the game ends should choice $\ell$ be selected by the bidder $\xi$. For a typical iterative auction, one generally has $\ell=2$ with the two sets corresponding to types above and below some value, respectively. Or, for a single-action auction, the $W_{\ell}$’s are generally singletons, with one per type in $V_{i}$.
We show that for any implementable outcome $(\tilde{x},\tilde{t})$ of the auction, one can always find a menu rule that is “informative”—the set of possible outcomes differs across partition selections [^24] This notion of informativeness is very similar to the pruned condition from Akbarpour and Li (2020).—that also implements the same outcome. So, without loss of generality, we restrict menu rules in this way and then describe an auction equilibrium by $(\tilde{x},\tilde{t},\mu,\xi_{0})$. (See Lemma A.5 in the Appendix for a more formal treatment.) We refer to $(\tilde{x},\tilde{t},\mu,\xi_{0})$ as an auction when convenient. As is always the case with direct mechanisms, the auction encompasses both the game form and the equilibrium, i.e., by appealing to the revelation principle we have implicitly selected the equilibrium already.
Finally, as notation for later sections, for a set $V=V_{1}\times V_{2}\times\cdots\times V_{N}$ of valuation profiles, define $\overline{{V}}_{i}=\max\left\{v_{i}:v_{i}\in V_{i}\right\}$ to be the maximum possible value of bidder $i$; $\overline{{V}}_{-i}$ similarly to be the maximum possible value of bidders $j\neq i$; and $\overline{{V}}=\max_{i}\left\{\overline{{V}}_{i}\right\}$. We define $\underline{{V}}_{i},\underline{{V}}_{-i}$, and $\underline{{V}}$ as the corresponding values for minima instead of maxima.
3 Dutch Auctions
Define the bidding rule [^25] See Lemma B.1 in the Appendix for proof that this is the correct form. The difference in transfer function by bidder is because of tie-breaking.
$b^{1}_{i}(v^{m})=v^{m}-\sum_{k:v^{k}<v^{m}}(v^{k+1}-v^{k})\frac{\left(F(v^{k})\right)^{i-1}\left(F(v^{k-1})\right)^{N-i-1}}{\left(F(v^{m})\right)^{i-1}\left(F(v^{m-1})\right)^{N-i-1}}.$
Then, the Dutch auction is defined as the auction which begins by offering each bidder $i$ the item at $b^{1}_{i}(v^{M})$, and then if no bidder chooses to buy the item at that price, the item is offered for $b^{1}_{i}(v^{M-1})$ and so on until either a bidder has chosen to buy the item or the price to be offered drops below $b^{1}_{i}(\rho^{*})$, where $\rho^{*}$ denotes an optimal reserve price. We consider only orderly auctions and therefore, at each price level, bidders are offered the opportunity to buy the item in priority order. Formally:
Definition 3.1.
The Dutch auction with reserve price $\rho^{*}$ is defined by the optimal allocation rule $\tilde{x}^{*}$, first-price transfer rule $\tilde{t}^{1}=\tilde{x}^{*}\cdot b^{1}$, initial player $\left(\cdot,\xi_{0}\right)=\max_{\rhd}\left\{\left(0,i\right)\right\}$, and menu
$\mu^{D}(V,\xi)$ =$=\left\{\left(W_{\mathsf{L}},\vec{\xi}_{\mathsf{L}}\right),\left(W_{\mathsf{H}},\vec{\xi}_{\mathsf{H}}\right)\right\},$
where $W_{\mathsf{H}}=\left\{\overline{{V}}_{\xi}\right\},W_{\mathsf{L}}=V_{\xi}\backslash\left\{\overline{{V}}_{\xi}\right\},\vec{\xi}_{\mathsf{H}}=\emptyset,\text{ and }$
$\vec{\xi}_{\mathsf{L}}=\begin{cases}\left(\cdot,\tilde{\xi}\right)=\max_{\rhd}\left\{\left(\overline{{V}}_{i},i\right):i\neq\xi\right\}&\exists i\neq\xi\text{ such that }\left|V_{i}\right|>1\text{ and }\overline{{V}}_{i}\geq\rho^{*}\\ \emptyset&\text{otherwise}\end{cases}.$
In the following subsections, we explore how the Dutch auction is uniquely suited to preventing shill bidding.
3.1 Strongly Shill-Proof Auctions
Lemma 3.2 (Pay-as-bid).
If an optimal auction $(\tilde{x}^{*},\tilde{t},\mu,\xi_{0})$ is strongly shill-proof, then it must be a pay-as-bid auction. Formally, for all $\xi,v_{\xi}$ and $v_{-\xi},v_{-\xi}^{\prime}$,
$\tilde{x}^{*}_{\xi}\left(v_{\xi},v_{-\xi}\right)=\tilde{x}^{*}_{\xi}\left(v_{\xi},v_{-\xi}^{\prime}\right)\implies\tilde{t}_{\xi}\left(v_{\xi},v_{-\xi}\right)=\tilde{t}_{\xi}\left(v_{\xi},v_{-\xi}^{\prime}\right).$
Observe that Lemma 3.2 (and Theorem 3.4) holds even if we relax the assumption on shill bidders’ information sets because an auction being strongly shill-proof means that shill bidders want to report 00 even if they know the precise valuations of other bidders ex-ante. To see why Lemma 3.2 is true, consider the case where $R=\left\{\xi\right\}$. Then, the shill bidders will report whichever values maximize $\tilde{t}_{\xi}$ and so $\tilde{t}_{\xi}$ must be constant across all outcomes with the same allocation.
To state our main result, we consider public auctions:

### v3

@@ 1.4 Outline of the Paper
In Section 2, we present our formal model of shill bidding in auctions. Section 3 studies strong shill-proofness and shows the ways in which Dutch auctions are uniquely optimal at disincentivizing shill bidding. Section 4 explores which formats are weakly shill-proof and provides a trilemma result. Section 5 generalizes the model to a setting with affiliation and interdependent values, proves that shill-proofness admits, and ordering with respect to affiliation and interdependent values, and generalizes our results. In Section 6, we conclude the paper by discussing extensions.
@@ 2 Model
In this paper, we consider extensive-form, single item auctions. An extensive-form game $G$ is a tuple of possible histories $H$, and, for each history $h\in H$, functions mapping $h$ to: (i) a player taking an action, $P(h)$; (ii) a set of possible actions, $A(h)$; (iii) an information set, $\mathcal{I}(h)$; and (iv) the most recent action taken, $\mathcal{A}(h)$. As further notation, we denote the starting history of the game by $h_{\emptyset}$ and the set of terminal histories as $Z$; we say $h^{\prime}\prec h$ if $h^{\prime}$ precedes $h$, i.e., there exists a sequence of actions that lead from $h^{\prime}$ to $h$.
We restrict attention to single-item auctions, which means that for every terminal history $z\in Z$, we can associate an allocation and transfer vector: $z=(x,t)$, with $\sum_{i=1}^{N}x_{i}\leq 1$ and $x_{i}\in\left\{0,1\right\}$ for all $i$. Abusing notation slightly, we use $x(z),t(z)$ to mean the vectors $(x,t)$ associated with the terminal history $z$. We also assume perfect recall and finite depth. (Definition E.1 in the Supplemental Appendix gives a formal, thorough, and standard definition of extensive-form games.)
@@ 2.1 Bidders – Real and Shill
[…]
Note that (i) whether a given bidder is real and (ii) the bidders’ values for the item are not built into the extensive-form game $G$. Instead, bidders’ strategies are a function of their types. Like most papers in the extensive-form auction literature, we study games with a finite type space because defining auctions with a continuum of types requires defining a general class of continuous-time games. To the authors’ knowledge, there is no theory of continuous-time games that rivals the generality and flexibility of extensive-form games.
@@
Real bidders have no ex-ante information about who else is a real bidder; they only know that each other bidder is real with probability pp. [^16] Unlike the assumption that $N$ is fixed, this assumption is an economically substantive one: The only way for shill bidders to manipulate the outcome of the auction is for shill bidders to take actions in the auction. If the bidders instead knew who were the real bidders, shill bidders could have an incentive to appear as real bidders in order to increase perceived competition. However, throughout the course of the auction, they can update their beliefs about which bidders are real and adjust their actions accordingly. We assume that shill bidders know the set of shill bidders and observe all previous actions taken. Formally, for any history $h$, if $P(h)\in S$, then $\mathcal{I}(h)=\left\{h\right\}$. This assumption rules out games with simultaneity (including static games) from the perspective of the shill bidders, but not real bidders. [^17] In extensive form games, simultaneity is modeled as the information set of a player having multiple elements. Without cryptography or other unmodeled technologies, we view it as reasonable to assume that while the auction may appear simultaneous to the real bidders, actions are taking place sequentially and the seller can observe those actions.
Our equilibrium concept is pure-strategy Perfect Bayesian Equilibrium; a formal definition of the auction equilibrium $(G,\sigma)$ can be found in Definition A.1. We write $\sigma(\theta;R)$ for the strategy profile when the value profile is $\theta$ and the realized set of real bidders is $R$. Perfect Bayesian equilibria are defined without consideration of group deviations by the shill bidders. However, with this setup, it is without loss to focus only on individual shill bidder’s incentives as opposed to any group actions because all the shill bidders have the same objective function and information available to them.
@@ 2.2 Auction Environment
Throughout the paper, we focus on auction equilibria that are ex-post monotone and individually rational: The auction equilibrium $(G,\sigma)$ is monotone if, for all $i$, $j$, $\theta_{-j}$, and $\theta_{j}>\theta_{j}^{\prime}$, $\left[t_{i}\left(\sigma(\theta;B)\right)>0\implies t_{i}\left(\sigma(\theta;B)\right)\geq t_{i}\left(\sigma(\theta_{j}^{\prime},\theta_{-j};B)\right)\right],$ and is individually rational (IR) if, for all $v$ and $i\in B$, $x_{i}(\sigma(\theta;B))v_{i}-t_{i}(\sigma(\theta;B))\geq 0.$ The monotonicity condition on transfers is satisfied by all standard single-item auction formats such as English auctions, Vickrey auctions, Dutch auctions, and sealed first-price auctions. Moreover, note that we impose the monotonicity condition primarily for convenience; all our results continue to hold if we instead just assume that total transfers are weakly higher whenever any bidder reports a value higher than 0. The ex-post IR condition rules out all-pay auctions and ensures that only the winner pays the seller. [^18] Unlike with continuous types, with discrete types there is positive probability of multiple bidders having the same highest value and so an optimal or efficient auction make randomize between bidders. For many of our results, we will assume Definition 3.2 to rule out this case. We also make this assumption primarily for convenience, versions of all our main results would still hold if we were to relax to ex-interim IR.
2.3 Shill-Proofness
Next, we define our key shill-proofness desiderata. We are interested in auction equilibria in which shill bidders do not shill. Formally, this corresponds to requiring that shill bidders always act like real bidders who have value 0 for the item—since real bidders who have value 0 will never enter non-trivial bids in equilibrium, requiring shill bidders to have the same actions in equilibrium in effect means that shilling does not occur.
Definition 2.1.
An auction equilibrium $(G,\sigma)$ is weakly shill-proof if $\sigma$ is invariant to the realization of $S$, i.e., for all $\theta$ and $S,S^{\prime}\subseteq\left\{i:\theta_{i}=0\right\}$: $\sigma\left(\theta;B\backslash S\right)=\sigma\left(\theta;B\backslash S^{\prime}\right)$.
We can also strengthen our no-shilling criterion from equilibrium to ex-post strategy:
Definition 2.2.
An auction equilibrium $(G,\sigma)$ is strongly shill-proof if it is weakly shill-proof and $\sigma$ is an ex-post strategy profile for shill bidders, i.e., for all $\sigma^{\prime}$, $S$, and $\theta_{-S}$,
$\sum_{j\in R}t_{j}(\sigma(0,\theta_{-S};R))\geq\sum_{j\in R}t_{j}\left(\sigma_{S}^{\prime},\sigma_{-S}(0,\theta_{-S};R)\right).$
Strong shill-proofness is obviously preferable to weak shill-proofness (all else equal), especially if there are concerns about a seller acquiring information about real bidders’ valuations beyond what is encoded by the prior. As we will see, however, the design space of weak shill-proof auctions is meaningfully larger than that of strong shill-proof auctions.
2.4 Revelation Principle
In order to make progress in understanding shill-proof auction formats, the following revelation principle will be helpful: for every auction equilibrium $\left(G,\sigma\right)$, there exists a direct auction that can be summarized by a direct allocation rule $\tilde{x}:\vartheta^{N}\to[0,1]^{N}$, a direct transfer rule $\tilde{t}:\vartheta^{N}\to\mathbb{R}^{N}$, a menu rule
$\mu:\mathcal{P}\left(\vartheta^{N}\right)\times B\to\bigcup_{L=2}^{M}\left(\left\{\mathcal{T}\in\Pi(\vartheta)\mid\left|\mathcal{T}\right|=L\right\}\times B^{L}\right)$
where $\Pi(\vartheta)$ is the set of all possible partitions over the type space, and a starting player $\xi_{0}\in B$. [^19] This revelation principle is similar to those found in, for example, Ashlagi and Gonczarowski (2018); Mackenzie (2020); Mackenzie and Zhou (2022); Pycia and Troyan (2023). The first input to the menu rule $\mu$ is a set $\Theta$ of valuation profiles of the form $\Theta=\Theta_{1}\times\Theta_{2}\times\cdots\times\Theta_{N}$ with $\Theta_{i}\subseteq\vartheta$ for all ii—intuitively, the valuation profiles that are, in equilibrium, consistent with a particular history. The second input is a player $\xi$ who is to move next. The output of the rule is a collection $\left\{\left(W_{\ell},\vec{\xi}_{\ell}\right)\right\}_{\ell\in\left\{1,\ldots,L\right\}}$, where the $W_{\ell}$’s are a partition of $\Theta_{\xi}$ (the equilibrium strategy $\sigma$ determines the partition; player $\xi$ will truthfully choose a subset based on her valuation) and $\vec{\xi}_{\ell}$ indicates the next player to move should player $\xi$ choose $W_{\ell}$. Under $\sigma$, the player $\xi$ will always select the partition $W_{\ell}$ such that $\theta_{\xi}\in W_{\ell}$. If $\vec{\xi}_{\ell}=\emptyset$, then the game ends should choice $\ell$ be selected by the bidder $\xi$. For a typical iterative auction, one generally has $L=2$ with the two sets corresponding to types above and below some value, respectively. Or, for a single-action auction, the $W_{\ell}$’s are generally singletons, with one per type in $\Theta_{i}$.
[…]
Finally, as notation for later sections, for a set $\Theta$ of valuation profiles, define $\overline{{\Theta}}_{i}=\max\left\{\theta_{i}:\theta_{i}\in\Theta_{i}\right\}$ to be the maximum possible value of bidder $i$; $\overline{{\Theta}}_{-i}$ to be the maximum possible value of bidders $j\neq i$; and $\overline{{\Theta}}=\max_{i}\left\{\overline{{\Theta}}_{i}\right\}$. We define $\underline{{\Theta}}_{i},\underline{{\Theta}}_{-i}$, and $\underline{{\Theta}}$ as the corresponding values for minima instead of maxima.
3 Strongly Shill-Proof Auctions
3.1 Direct Mechanisms
Lemma 3.1 (Pay-as-bid).
If an auction $(\tilde{x},\tilde{t},\mu,\xi_{0})$ is strongly shill-proof, then it must be a pay-as-bid auction. Formally, for all $\xi,\theta_{\xi}$, and $\theta_{-\xi},\theta_{-\xi}^{\prime}$,
$\tilde{x}_{\xi}\left(\theta_{\xi},\theta_{-\xi}\right)=\tilde{x}_{\xi}\left(\theta_{\xi},\theta_{-\xi}^{\prime}\right)\implies\tilde{t}_{\xi}\left(\theta_{\xi},\theta_{-\xi}\right)=\tilde{t}_{\xi}\left(\theta_{\xi},\theta_{-\xi}^{\prime}\right).$
The proof of Lemma 3.1 and all other results can be found in the Appendix. Observe that Lemma 3.1 holds (as does Theorem 3.6) even if we relax the assumption on shill bidders’ information sets because strong shill-proofness means that shill bidders want to report 0 even if they know the precise valuations of other bidders ex-ante. To see why Lemma 3.1 is true, consider the case where $R=\left\{\xi\right\}$. Then, the shill bidders will report whichever values maximize $\tilde{t}_{\xi}$ and so $\tilde{t}_{\xi}$ must be constant across all outcomes with the same allocation.
@@ 3.1 Direct Mechanisms
Recall that with continuous types, two bidders have the same type with probability 0 and the optimal allocation rule is uniquely defined up to measure-zero events. However, since types are discrete, the probability of ties is non-zero and so we have to define a tie-breaking rule. We adopt the notion of orderliness introduced by Akbarpour and Li (2020): there exists a fixed priority order—independent of values—over which bidder wins an item if there is a tie. [^21] For example, if ties are broken lexicographically, then the auction is orderly. We clarify their definition by explicitly linking the priority order to the ex-post allocation rule.
@@
    $\theta_{i}>\theta_{j}\implies(\theta_{i},i)\rhd(\theta_{j},j)$;
@@
    if there exists $m$ such that $(\theta^{m},i)\rhd(\theta^{m},j)$, then for all $k$, $(\theta^{k},i)\rhd(\theta^{k},j)$; and
  3. (iii)
    for all $\theta$, if $\tilde{x}_{i}(\theta)\in\left\{0,1\right\}$ and $(\theta_{i},i)\rhd(\theta_{j},j)$, then $\tilde{x}_{i}(\theta)>\tilde{x}_{j}(\theta)$.
We can now define the orderly allocation rule for any auction with reserve type $\theta^{\rho}$ as
$\tilde{x}^{\rho}_{i}(\theta)=\mathds{1}\left\{\theta_{i}\geq\theta^{\rho},\left(\theta_{i},i\right)=\max_{\rhd}\left\{\left(\theta_{j},j\right)\right\}_{j\in B}\right\}.$
Then, the orderly, efficient allocation rule is $\tilde{x}^{E}=\tilde{x}^{1}$ and the orderly, optimal allocation rule is $\tilde{x}^{*}=\tilde{x}^{\rho^{*}}$ for some $\rho^{*}$.
We can now explicitly define the revenue-maximizing pay-as-bid bidding functions. Note that we need to select among multiple IC bidding functions because the discrete type space implies that the IC constraints need not bind with equality. Using Lemma B.2 from the Appendix, the bidding function for bidder $i$ is given by
$b^{1}_{i}(\theta^{m})=\theta^{m}-\sum_{k:\theta^{k}<\theta^{m}}(\theta^{k+1}-\theta^{k})\frac{\left(F(\theta^{k})\right)^{i-1}\left(F(\theta^{k-1})\right)^{N-i-1}}{\left(F(\theta^{m})\right)^{i-1}\left(F(\theta^{m-1})\right)^{N-i-1}}.$
The transfer rule is then $\tilde{t}^{E}=b^{1}\cdot\tilde{x}^{E}$ and $\tilde{t}^{*}=b^{1}\cdot\tilde{x}^{*}$ for the efficient and optimal auctions, respectively. Observe that the asymmetry in the bidding and transfer functions arises purely from the discrete types and the tie-breaking rule, not from the definition of shill-proofness.
3.2 Indirect Mechanisms
Now that we have pinned down the direct mechanism, we turn to the indirect implementation. The information structure available to real bidders matters for which mechanisms are strongly shill-proof. If there is no information leakage between bidders, then the classic first-price auction is strongly shill-proof. However, if any bidder can learn any outcome-relevant information about another bidder’s actions, then the first-price auction is not strongly shill-proof. In Section 3.3, we return to the primary setting studied in this work, public auctions.
To expand, the single-action, first-price auction where each bidder sequentially makes a bid will be strongly shill-proof when the game is static, i.e., when each bidder takes her single action with no information about what other bidders have done. However, we can show that the single-action, first-price auction is strongly shill-proof only in static games—if there is ever a chance that a bidder learns some outcome-informative information about what actions previous bidders have taken, then the auction is no longer strongly shill-proof. Formally, in an orderly auction, for some $i,j$, and $\theta$ with $(\theta^{M},j)\rhd(\theta_{i},i)\rhd(\theta^{\rho^{*}},j)$, let $h_{\theta}$ be the (unique) history reached in equilibrium by $\sigma(\theta;B)$ where $P(h)=j$ and let $h_{0}$ be the corresponding history reached by types $(0,\theta_{-i})$. We say that an information structure is strictly more informative than a static game if there exist $i$, $j$, and $\theta$ such that $\mathcal{I}(h_{\theta})\neq\mathcal{I}(h_{0})$ and $\tilde{x}_{j}(\theta)=1$. Note that if there exists such a type vector, then there must exist a type vector where $\tilde{x}_{j}(\theta)=0$.
Proposition 3.3.
Under a static information structure, the single-action, first-price auction is strongly shill-proof. However, for any information structure strictly more informative than a static game, the single-action, first price auction is not strongly shill-proof.
A core intuition underlying Proposition 3.3 is that the stronger a bidder perceives her competition, the less she shades her bid in a first-price auction. So, if a shill bidder $i$ is able to distinguish that she has a value that could change the outcome of the auction to bidder $j$, then there must be some value $\theta_{j}$ under bidder $j$ perceives the auction as more competitive; this is occurs if and only if there is outcome-relevant information leaked from bidder $i$ to $j$. This will increase revenue if $j$ wins and so the first-price auction is not strongly shill-proof.
As noted in Section 2, implementing no information leakage (or, equivalently, simultaneity) between bidders is often an unrealistic assumption. Even with some technology that might deter information leakage in the actual mechanism, information leakage through side channels is inevitable in many settings. So, we take Proposition 3.3 as evidence therefore that the first-price auction is, in many settings, potentially subject to manipulation by shill bidders.
3.3 Dutch Auctions
Now that we have shown that the extensive form of an auction affects whether or not it is strongly shill-proof, we turn to studying which extensive forms are always strongly shill-proof. We show that Dutch auctions are strongly shill-proof, no matter the information structure, and show that that there exists a natural information structure—the public setting—such that Dutch auctions are uniquely strongly shill-proof.
To begin, we formally define the public information structure: